Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $127,400 of manufacturing overhead for an estimated allocation base of $91,000 direct material dollars to be used in production. The company has provided the following data for the just completed year: Purchase of raw materials $ 139,000 Direct labor cost $ 87,000 Manufacturing overhead costs: Indirect labor $ 126,300 Property taxes $ 9,000 Depreciation of equipment $ 20,000 Maintenance $ 13,000 Insurance $ 9,800 Rent, building $ 36,000 Beginning Ending Raw Materials $ 27,000 $ 16,000 Work in Process $ 45,000 $ 37,000 Finished Goods $ 75,000 $ 61,000 Required: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhead for the year. 3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials. 4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer. 5. Assume that the $37,000 ending balance in Work in Process includes $8,700 of direct materials. Given this assumption, supply the information missing below:

Assume that the $37,000 ending balance in Work in Process includes $8,700 of direct materials. Given this assumption, supply the information missing below:

Direct materials $8,700
Direct labor
Manufacturing overhead
Work in process inventory $37,000
Answer #1
1
Estimated manufacturing overhead 127400
Divide by estimated direct materials 91000
Predetermined overhead rate 140%
2
Indirect labor 126300
Property 9000
Depreciation of equipment 20000
Maintenance 13000
Insurance 9800
Rent, building 36000
Total Manufacturing overhead costs 214100
Total Actual Manufacturing overhead costs 214100
Less: Manufacturing overhead applied to work in process 210000 =150000*140%
Underapplied overhead 4100
3
Schedule of Cost of Goods Manufactured
Direct materials:
Beginning Raw materials inventory 27000
Add: Purchases of raw materials 139000
Total raw materials available 166000
Less: Ending Raw materials inventory 16000
Raw materials used in production 150000
Direct labor cost 87000
Manufacturing overhead applied to work in process 210000 =150000*140%
Total manufacturing costs 447000
Add: Beginning Work in process inventory 45000
492000
Less: Ending Work in process inventory 37000
Cost of goods manufactured 455000
4
Finished goods inventory, beginning 75000
Add: Cost of goods manufactured 455000
Cost of Goods available for sale 530000
Less: Finished goods inventory, ending 61000
Unadjusted Cost of goods sold 469000
5
Direct materials 8700
Direct labor 16120 =37000-8700-12180
Manufacturing overhead 12180 =8700*140%
Work in process inventory 37000